# Modeling Large Nonstationary Spatial Data with the Full-Scale Basis Graphical Lasso

### v1

@@ 2.2.1 Estimation for the Full-Scale Basis Graphical Lasso (FSBGL)
The hat matrix is the optimal predictor of the field given the data, and its trace, which lies between 00 and $J+1$, is a measure of the influence of $Q$. If the trace is small, that means that $Q$ is large, i.e. the estimates of $\beta$ have small variance, and if the trace is large, then our estimates of $\beta$ have large variance. For this reason, keeping $\text{tr}\left(\hat{H}\right)$ small prevents overfitting.
3 Illustration
@@ 3.1 Data
The National Center for Atmospheric Research (NCAR) Whole Atmosphere Community Climate Model with thermosphere and ionosphere extension (WACCM-X) is a high-top Earth system atmosphere model that extends from the surface to low-Earth orbit (LEO) altitudes ($\sim$500 km) ([waccmx]). Recently, WACCM-X was run at a high-resolution configuration to model intricate structures of the thermosphere influenced by atomospheric wave forcing due to highly variable tropospheric weather ([liu.gravity.waves]). Variability of the lower thermospheric temperature structures (80-150km) is not well represented in the standard atmosphere models, but it is critical to modeling aerodynamic forces experienced by vehicles and objects during reentry ([leonardetal]). Thermospheric temperature fields derived from the data used in [liu.gravity.waves] are used for illustrations of the method.
The data we examine consist of 41 replicate samples of neutral temperature fields at a pressure level corresponding roughly to 130-135 km altitude. The data correspond to simulation outputs with a 6 hour time resolution from 0:00 UT on January 13 to 24 UT January 22 under solar minimum conditions. These data are available on a $0.25^{\text{o}}\times 0.25^{\text{o}}$ grid, but are downsampled to $2^{\text{o}}\times 2^{\text{o}}$ for computational concerns. This also allows us to test the model’s ability to predict spatial scales smaller contained in training data, which are referred to as subgrid in this study. The model is fit to data from January 13 to January 20 (32 samples) and tested on the remaining data (9 samples). The spatial coordinate system is rotated with the motion of the Sun to focus on spatial variability that is synchronized with solar heating. The center of the figures shown in this paper thus corresponds to local noon, and the edges to local midnight. The empirical mean of these data, as well as three fields with the mean subtracted, computed on the $2^{\text{o}}\times 2^{\text{o}}$ grid, are shown in Figure 1. The dimension of these data, containing over 16000 spatial locations for each field, and the complicated structure seen in Figure 1 challenge existing modeling frameworks.
@@
In the next subsection we detail the FSBGL modeling choices and in-sample model fit, followed by a detailed comparison against competing alternatives.
@@ 3.2.1 $Z_{1}(x)$
For the low-rank portion of the model, we choose to write the field in a needlet expansion. Needlets are a multiresolution frame for data on $\mathbb{S}^{2}$ first introduced in [needletintro], and additional details of their construction are given in [cmbneeds] and [RFSphere]. Letting $Y_{lm}(x)$ be the spherical harmonics, the needlet $\psi_{jk}(x)$ is given by
@@
Here, $j$ is a spatial resolution parameter, $k$ controls location, $B$ controls the extend of the needlets in harmonic space (and here will be set to 2), $\{\lambda_{jk},\xi_{jk}\}$ form an exact cubature rule for polynomials of degree $\lceil B^{j+1}\rceil$, and $b(t)$ is a $C^{\infty}$ function with compact support on the interval $[B^{-1},B]$ such that $\forall\eta>1,\sum_{j}b(\frac{\eta}{B^{j}})=1$. For such $b(t)$, the needlets are compactly supported in the harmonic domain and additionally decay faster than any rational function away from $\xi_{jk}$, meaning that they have excellent localization properties even though they do not have local support ([RFSphere]).
In addition to providing localization in physical and harmonic space, they have several other important analytic properties. Along the first spherical harmonic $Y_{00}=\frac{1}{\sqrt{4\pi}}=\psi_{-1,0}$ they form a Parseval-tight frame of $\mathcal{L}^{2}\left(\mathbb{S}^{2}\right)$ ([needletintro]). This means that for all $f(x)\in\mathcal{L}^{2}\left(\mathbb{S}^{2}\right)$,
@@
$\left|\left|f\right|\right|_{2}^{2}$ $=\sum_{j,k}\left|\beta_{jk}\right|^{2}$ (15)
@@
so they have many of the same reconstruction properties as orthonormal bases while maintaining additional flexibility ([Intro_frames]). Additionally, the coefficients are asymptotically decorrelated as $j\to\infty$, and due to their local support this holds even if there are large swathes of missing data ([RFSphere]).
We will use the MATLAB package NeedMat to calculate the needlets ([fan]).
@@ 3.2.2 $Z_{2}(x)$
We consider a suite of common compactly supported options for the small-scale process $Z_{2}(x)$. The first class of correlation functions we will entertain are the Gaspari-Cohn correlation functions first described in [gasparicohn]. This class of correlation functions is generated by a self-convolution of a family of piecewise-linear functions over the domain of interest. This family of piecewise-linear functions has the parametrization
@@
where $a$ is a shape parameter and $c$ a scale parameter. This family exhibits negative correlation for $a=-0.1$, and to distinguish this from the other correlation models we are choosing to fix $a$ to this value (See [gasparicohn] Figure 7). The Gaspari-Cohn correlation function is then given by
@@
It is important to note that the Gaspari-Cohn correlation function is not positive-definite on the sphere ([gneitingpdsphere]), which means that we must use the chordal distance $d=2\sin\left(\frac{\alpha}{2}\right)$, where $\alpha=\arccos{\left\langle x_{1},x_{2}\right\rangle}$ is the angle between the position vectors of the two points. While this can lead to unusual correlation structures over large distances, since the model will be compactly supported and describe local variation we are not concerned with the warping of the surface that this will introduce.
The second class of covariance models that we consider is a mixture of Wendland covariance functions ([wendlandrbf]). The Wendland covariance is a radial basis function with compact support outside a cutoff radius $\theta$ of the form
@@
Since they are positive definite functions on the sphere ([gneitingpdsphere]), we will use the angular distance measure $d=\arccos{\left\langle x_{1},x_{2}\right\rangle}$. We propose to model the covariance of the small-scale process using the mixture of Wendland radial basis functions as
@@
Lastly, we will investigate a tapered Matérn correlation model ([furrercovtapering]). This model is given by
$C(d)=C_{\nu}\left(\frac{d}{r}\right)C_{\theta}(d)$ (20)
where $K_{\nu}$ is a modified Bessel function of the second kind. Fields with a Matérn correlation have $m$ mean-square derivatives if and only if $\nu>m$ ([steinbook]), which allows us to fit the smoothness of the data directly, unlike the other models examined here. In addition to optimizing over $\nu$ and $r$, we will also fit the range parameter $\theta$ of the taper function, which we have chosen to be a Wendland radial basis function of the form in (18). Again, this model is positive-definite on the sphere, so we will let $d$ be the angular distance.
@@ 3.3.1 Model Selection
Model cAIC $\text{tr}\left(\hat{H}\right)$
Gaspari-Cohn $a=-0.1$ $1.99\times 10^{5}$ $48.31$
@@
Table 1: Conditional AIC for the models of $Z_{2}$ in Section 3.2.2.
The fitted parameters for the four models tested are given in Table 2. The most striking result is the significant variation in the value of the fitted nugget variance $\tau^{2}$. While the tapered Matérn model has a fitted nugget variance of just $0.15$, the fitted variance for other models is at least 100 times larger, and in the case of the Gaspari-Cohn model it is about 650 times larger. This variation is likely caused by differences in how well each model can handle the underlying smoothness of the data: a stationary field whose correlation function has $2k$ derivatives at the origin will be $k$ times mean-square differentiable ([steinbook]) and both the Wendland and Gaspari-Cohn correlation functions have derivatives at the origin, implying that the underlying field $Z_{2}$ has some degree of smoothness. However, this smoothness is not supported by the data, and so to compensate for that the nugget variance is driven higher in an attempt to match the observed smoothness. Since for the tapered Matérn model we are effectively fitting the number of derivatives of the field the model is able to compensate for this directly, and does, selecting $\nu\approx 0.63$, corresponding to a mean-square continuous but not differentiable $Z_{2}$. This means that the nugget effect is not responsible for matching the underlying smoothness of the field all on its own, and this leads to a much smaller variance being fit.
@@ 3.4.1 LatticeKrig
The LatticeKrig model introduced in [nychka] is a multiresolution model in which the field is described via Wendland radial basis functions and the coefficients are modeled as a Gaussian Markov random field with node locations at the centers of the radial basis functions. By describing the coefficients with a spatial autoregression, $Q$ is guaranteed to be a sparse matrix, and the compact support of the basis functions introduces sparsity in the basis matrix $\Psi$ as well. This sparsity can be leveraged to allow the model to be evaluated extremely quickly, making likelihood computations fast and minimizing the computational load of model fitting. In addition, the structure allows the LatticeKrig model to adapt to a wide variety of spatial covariance functions ([nychka]). In a spherical geometry the LatticeKrig model has a similar number of basis functions to the needlet frame at each level of resolution, so we assume three levels of resolution as we did for the FSBGL.
LatticeKrig is a stationary model and due to its parametrization. This means that it performs somewhat more poorly in this context than the methods we have discussed, which are able to infer some level of nonstationarity through the data to inform the structure of $Q$. The negative log-likelihood of LatticeKrig on the test set to is approximately $1.36\times 10^{5}$, far higher than any of the other models (see Table 3). As will be discussed in more detail later, it is also the poorest predictor of subgrid-scale data as it is not able to leverage any knowledge of nonstationarity.
@@ 3.4.2 Basis Graphical Lasso
The principal difference between our construction and that of the basis graphical lasso lies in the process $Z_{2}$. While the basis graphical lasso assumes that $Z_{2}=0$, our extension has generalized this to a stationary, compactly supported covariance. This allows much greater flexibility in the modeling of spatial datasets with multiscale structure, since, as noted in [steinlowrank], low-rank models like the basis graphical lasso have trouble recreating small-scale variability. This is demonstrated clearly in Figure 6(a), where the basis graphical lasso is able to capture large scale variation reasonably well but is not able to capture the local variation, specially in comparison to the FSBGL and full-scale approximations shown in Figures 6(b) and 6(c). Additionally, when applied to the test dataset, while the BGL is clearly superior to LatticeKrig, it is outclassed by the other models by a similar amount.
(a) Two draws from the distribution fit with the basis graphical lasso
(b) Two draws from the distribution fit with the full-scale approximation
(c) Two draws from the distribution fit with the FSBGL-TM model
Figure 6: Draws from the fitted models for the BGL, full-scale approximations, and FSBGL-TM. The Latter two show much more structure than the BGL, which does not contain any way of modeling small-scale correlation structure that is not included in the basis.
@@ 3.4.3 Full-Scale Approximation
This model is motivated by the full-scale approximations (FSA) of [steinfullscale] and [sangfullscale], so we wish to compare the results of applying both methods. Due to the limitations of the model proposed by [steinfullscale], where the entire covariance matrix $Q^{-1}$ is estimated, severely limiting the number of basis functions, we will be following the specification laid out in [sangfullscale]. There, the low-rank model consists of EOFs calculated from the same 32 realizations as the other algorithms presented, and we selected the EOFs that explained $70$ percent of the variance seen in the data. This value was chosen by examination of the fitted low-rank portion of the FSBGL, where the low-rank portion explained less than $70\%$ of the variance at $90\%$ of spatial locations. Since the EOFs are generated from the data, we expect to see that they perform the best on the in-sample data. This performance is demonstrated in Figure 6(b), with the FSA matching the statistical behavior of the data reasonably well. Additionally, calculating the EOFs is fast, and their orthonormality simplifies the computations of the parameters of $C_{2}$.

### v2

@@ 2.2.1 Estimation for the Full-Scale Basis Graphical Lasso (FSBGL)
[…]
3 Simulation Studies
In this section we consider a set of simulation studies to investigate whether the proposed FSBGL is able to recover the structure in $Q$, as well as accurately determine the parameters of the stationary residual process. We simulate random fields $Z(x)$ on $0,1]^{2}$ of the form
$Z(x)=\sum_{i=0}^{J}\beta_{i}\phi_{i}(x)+Z_{2}(x)+\epsilon(x)$
using basis functions $\frac{1}{2}\cos\left(2\pi mx\right)\cos(2\pi ny)$ for $m,n=0,1,..,10$, leading to 121 total basis functions.
Figure 1: Three different models of $Q$ used to test the ability of the FSBGL to recover structure in a known precision matrix in the simulation studies.
We consider multiple common types of precision matrix structure, and use a tapered Matérn covariance model for $Z_{2}(x)$ ([10]). The small-scale variation covariance is given by
$C(d)=C_{\nu}\left(\frac{d}{r}\right)C_{\theta}(d)$ (14)
where $K_{\nu}$ is a modified Bessel function of the second kind. Fields with a Matérn correlation have $m$ mean-square derivatives if and only if $\nu>m$ [35]. The taper function is a Wendland radial basis function [38]. We fit every parameter of the tapered Matérn model, along with a process variance and the logarithm of the nugget effect variance $\tau^{2}$, totaling 5 parameters.
We investigate the ability to recover the structure of $Q$ and the residual parameters in C. The focus is on three different models for $Q$, which we call the block-diagonal, hub and spoke, and random graph models. These models are shown in Figure 1. Results are averaged over 50 independent simulations of the random fields. The results for the block diagonal model are presented in Table 1. For the errors in the parameters we present the median absolute error in recovery over 50 simulations for each number of samples, and for the errors in $Q$ we calculate the relative error in the Frobenius norm $\|\hat{Q}-Q\|_{F}/\|Q\|_{F}$. In general we find that the model is able to recover the parameters of $Z_{2}$ with reasonable accuracy regardless of the number of samples, but the recovery of $Q$ is dependent on the number of samples.
Parameter Truth m=10 m=30 m=50 m=100
$Q$ block-diag 0.610 0.444 0.436 0.277
Missed non-zero (%) N/A 51.6 13.8 8.73 0
Missed zero (%) N/A 0.25 4.8 1.54 19.2
$\sigma^{2}$ 1 0.041 0.041 0.059 0.041
$\log_{10}(\tau^{2})$ -2 0.224 0.184 0.232 0.236
$\theta$ 0.3 0.027 0.019 0.043 0.035
$r$ 0.15 0.027 0.043 0.036 0.044
$\nu$ 0.5 0.043 0.073 0.062 0.060
$L(\hat{Q},\hat{p})/L(Q,p)$ 1 1.1 1.1 1.04 1.1
Table 1: Median error in recovery of residual parameters for varying number of independent replicates of the field using a block-diagonal model for $Q$, along with true values. All calculations are the median over 50 trials for a given number of independent replicates of the field. The error in the recovery of $Q$ is given by $\|\hat{Q}-Q\|_{F}/\|Q\|_{F}$ with $\|\cdot\|$ the Frobenius norm, and the error for the other parameters is the absolute error.
For the purposes of this study the errors for $Q$ are calculated for $\lambda=10^{k}$ for k=1,0.5,0,0.5,1k=-1,-0.5,0,0.5,1 and the reported values are the errors for the value of $\lambda$ that minimizes the errors. The optimal value of $\lambda$ decreases with the number of samples, as the sparsity penalty becomes less important with greater amounts of data. For small numbers of samples, then, the estimated $Q$ is typically diagonal, with additional structure recovered as the number of samples increases. The error in $Q$ is comparable to that achieved by the BGL in Table 1 of [20]. The Appendix contains results for the other precision matrix structures, whose results are analogous to the block-diagonal case. We also include visualizations of the estimated precision matrices, which suggest that the general precision structure is well-recovered once a couple dozen realizations are available.
4 Application to WACCM-X Neutral Temperature Fields
@@ 4.1 Data
The National Center for Atmospheric Research (NCAR) Whole Atmosphere Community Climate Model with thermosphere and ionosphere extension (WACCM-X) is a high-top Earth system atmosphere model that extends from the surface to low-Earth orbit (LEO) altitudes ($\sim$500 km) ([24]). Recently, WACCM-X was run at a high-resolution configuration to model intricate structures of the thermosphere influenced by atomospheric wave forcing due to highly variable tropospheric weather ([25]). Variability of the lower thermospheric temperature structures (80-150km) is not well represented in the standard atmosphere models, but it is critical to modeling aerodynamic forces experienced by vehicles and objects during reentry ([22]). Thermospheric temperature fields derived from the data used in [25] are used for illustrations of the method.
The data we examine consist of 41 replicate samples of neutral temperature fields at a pressure level corresponding roughly to 130-135 km altitude. The data correspond to simulation outputs with a 6 hour time resolution from 0:00 UT on January 13 to 24 UT January 22 under solar minimum conditions. These data are available on a $0.25^{\text{o}}\times 0.25^{\text{o}}$ grid, but are downsampled to $2^{\text{o}}\times 2^{\text{o}}$ to alleviate computational concerns. This also allows us to test the model’s ability to predict spatial scales smaller than those contained in the training data, which are referred to as subgrid in this study. The model is fit to data from January 13 to January 20 (32 samples) and tested on the remaining data (9 samples). The spatial coordinate system is rotated with the motion of the Sun to focus on spatial variability that is synchronized with solar heating. The center of the figures shown in this paper thus corresponds to local noon, and the edges to local midnight. The empirical mean of these data, as well as three fields with the mean subtracted, computed on the $2^{\text{o}}\times 2^{\text{o}}$ grid, are shown in Figure 2. Additionally, the data generally show little autocorrelation between samples, as shown in D. Dynamics in this altitude region is known to be dominated by the day-to-day variability of atmospheric waves originating from tropospheric weather. This allows us to treat the samples as independent in our analysis.
@@
The dimension of these data, containing over 16000 spatial locations for each field, and the complicated structure seen in Figure 2 challenge existing modeling frameworks. In the next subsection we detail the FSBGL modeling choices and in-sample model fit, followed by a detailed comparison against competing alternatives.
@@ 4.2.1 $Z_{1}(x)$
For the low-rank portion of the model, we choose to write the field in a needlet expansion. Needlets are a multiresolution frame for data on $\mathbb{S}^{2}$ first introduced in [30], and additional details of their construction are given in [28] and [27]. Letting $Y_{lm}(x)$ be the spherical harmonics, the needlet $\psi_{jk}(x)$ is given by
@@
Here, $j$ is a spatial resolution parameter, $k$ controls location, $B$ controls the extent of the needlets in harmonic space (and here will be set to 2), the weights and points $\{\lambda_{jk},\xi_{jk}\}$ form an exact cubature rule on $\mathbb{S}^{2}$ for polynomials of degree less than $B^{j+1}$, and $b(t)$ is a non-negative $C^{\infty}$ function with compact support on the interval $B^{-1},B]$ such that $\forall\eta>1,\sum_{j}b(\frac{\eta}{B^{j}})^{2}=1$. For such $b(t)$, the needlets are compactly supported in the harmonic domain and additionally decay faster than any rational function away from $\xi_{jk}$, meaning that they have excellent localization properties even though they do not have local support ([27]).
In addition to providing localization in physical and harmonic space, they have several other important analytic properties. Along with the first spherical harmonic $Y_{00}=\frac{1}{\sqrt{4\pi}}=\psi_{-1,0}$ they form a Parseval-tight frame of $\mathcal{L}^{2}\left(\mathbb{S}^{2}\right)$ ([30]). This means that for all $f(x)\in\mathcal{L}^{2}\left(\mathbb{S}^{2}\right)$,
@@
[…]
@@
so they have many of the same reconstruction properties as orthonormal bases while maintaining additional flexibility ([3]). Additionally, the coefficients are asymptotically decorrelated as $j\to\infty$, and due to their local support this holds even if there are large swathes of missing data ([27]). We use the MATLAB package NeedMat to calculate the needlets ([9]).
@@ 4.2.2 $Z_{2}(x)$
We consider a suite of common compactly supported options for the small-scale process $Z_{2}(x)$. The first class of correlation functions we will entertain are the Gaspari-Cohn correlation functions first described in [11]. This class of correlation functions is generated by a self-convolution of a family of piecewise-linear functions over the domain of interest. This family of piecewise-linear functions has the parametrization
@@
where $a$ is a shape parameter and $c$ a scale parameter. This family exhibits negative correlation for a=0.1a=-0.1, and to distinguish this from the other correlation models we are choosing to fix $a$ to this value (See [11] Figure 7). The Gaspari-Cohn correlation function is then given by
@@
It is important to note that the Gaspari-Cohn correlation function is not positive-definite on the sphere ([12]), which means that we must use the chordal distance $d=2\sin\left(\frac{\alpha}{2}\right)$, where $\alpha=\arccos{\left\langle x_{1},x_{2}\right\rangle}$ is the angle between the position vectors of the two points. While this can lead to unusual correlation structures over large distances, since the model will be compactly supported and describe local variation we are not concerned with the warping of the surface that this will introduce.
The second class of covariance models that we consider is a mixture of Wendland covariance functions ([38]). The Wendland covariance is a radial basis function with compact support outside a cutoff radius $\theta$ of the form
@@
Since they are positive definite functions on the sphere ([12]), we will use the angular distance measure $d=\arccos{\left\langle x_{1},x_{2}\right\rangle}$. We propose to model the covariance of the small-scale process using the mixture of Wendland radial basis functions as
@@
Lastly, we investigate the tapered Matérn correlation model discussed in Section 3. This model is also positive-definite on the sphere, so we will let $d$ be the angular distance.
@@ 4.3.1 Model Selection
[…]
Gaspari-Cohn a=0.1a=-0.1 $1.99\times 10^{5}$ $48.31$
@@
Table 2: Conditional AIC and trace of the hat matrix for the models of $Z_{2}$ in Section 4.2.2.
The fitted parameters for the four models tested are given in Table 3. The most striking result is the significant variation in the value of the fitted nugget variance $\tau^{2}$. While the tapered Matérn model has a fitted nugget variance of just $0.15$, the fitted variance for other models is at least 100 times larger, and in the case of the Gaspari-Cohn model it is about 650 times larger. This variation is likely caused by differences in how well each model can handle the underlying smoothness of the data: a stationary field whose correlation function has $2k$ derivatives at the origin will be $k$ times mean-square differentiable ([35]) and both the Wendland and Gaspari-Cohn correlation functions have derivatives at the origin, implying that the underlying field $Z_{2}$ has some degree of smoothness. However, this smoothness is not supported by the data, and so to compensate for that the nugget variance is driven higher in an attempt to match the observed smoothness. Since for the tapered Matérn model we are effectively fitting the number of derivatives of the field the model is able to compensate for this directly, and does, selecting $\nu\approx 0.63$, corresponding to a mean-square continuous but not differentiable $Z_{2}$. This means that the nugget effect is not responsible for matching the underlying smoothness of the field all on its own, and this leads to a much smaller variance being fit. This phenomenon is further investigated in E.
@@ 4.4.1 LatticeKrig
The LatticeKrig model introduced in [31] is a multiresolution model in which the field is described via Wendland radial basis functions and the coefficients are modeled as a Gaussian Markov random field with node locations at the centers of the radial basis functions. By describing the coefficients with a spatial autoregression, $Q$ is guaranteed to be a sparse matrix, and the compact support of the basis functions introduces sparsity in the basis matrix $\Psi$ as well. This sparsity can be leveraged to allow the model to be evaluated extremely quickly, making likelihood computations fast and minimizing the computational load of model fitting. In addition, the structure allows the LatticeKrig model to adapt to a wide variety of spatial covariance functions ([31]). In a spherical geometry the LatticeKrig model has a similar number of basis functions to the needlet frame at each level of resolution, so we assume three levels of resolution as we did for the FSBGL.
LatticeKrig is a stationary model and due to its parametrization cannot capture the full variability present in the data. This means that it performs somewhat more poorly in this context than the methods we have discussed, which are able to infer some level of nonstationarity through the data to inform the structure of $Q$. The negative log-likelihood of LatticeKrig on the test set to is approximately $1.36\times 10^{5}$, far higher than any of the other models (see Table 4). As will be discussed in more detail later, it is also the poorest predictor of subgrid-scale data as it is not able to leverage any knowledge of nonstationarity.
@@ 4.4.2 Basis Graphical Lasso
The principal difference between our construction and that of the basis graphical lasso lies in the process $Z_{2}$. While the basis graphical lasso assumes that $Z_{2}=0$, our extension has generalized this to a stationary, compactly supported covariance. This allows much greater flexibility in the modeling of spatial datasets with multiscale structure, since, as noted in [37], low-rank models like the basis graphical lasso have trouble recreating small-scale variability. Additionally, when applied to the test dataset, while the BGL is clearly superior to LatticeKrig, it is outclassed by the other models by a similar amount.
@@ 4.4.3 Full-Scale Approximation
This model is motivated by the full-scale approximations (FSA) of [36] and [33], so we wish to compare the results of applying both methods. Due to the limitations of the model proposed by [36], where the entire covariance matrix $Q^{-1}$ is estimated, severely limiting the number of basis functions, we will be following the specification laid out in [33]. There, the low-rank model consists of EOFs calculated from the same 32 realizations as the other algorithms presented, and we selected the EOFs that explained $70$ percent of the variance seen in the data. This value was chosen by examination of the fitted low-rank portion of the FSBGL, where the low-rank portion explained less than $70\%$ of the variance at $90\%$ of spatial locations. Since the EOFs are generated from the data, we expect to see that they perform the best on the in-sample data. Additionally, calculating the EOFs is fast, and their orthonormality simplifies the computations of the parameters of $C_{2}$.